1. The Green Company makes tables. Job 489 is to build 20 tables.

Job 489 requires $680 of direct materials and 30 hours of direct labor costing $530. Also use the following information about the company:

Total fixed manufacturing overhead cost estimated for all jobs during the whole year     $480,000                                   

Estimated direct labor hours for all jobs during the whole year                                                  60,000   

Estimated variable manufacturing overhead cost per direct labor hour                                          $3                                     

Calculate the following if the Green Company uses the traditional method of job order costing with one company-wide predetermined overhead rate using direct labor hours as the allocation base:

a. The predetermined overhead rate

b. The manufacturing overhead cost applied to job 489

c. The total cost of job 489

d. The cost per unit (per table)
Answer #1

a) Predetermine overhead rate = (480000/60000+3) = 11 per labor hour

b) Manufacturing overhead applied = 30*11 = 330

c) Total cost of Job 489

Direct material 680
Direct labor 530
Overhead 330
Total cost of Job 489 1540

d) The cost per unit = 1540/20 = $77